# Noise-augmented Chaotic Ising Machines for Combinatorial Optimization and Sampling

### v1

@@ IV.   Adaptive Parallel Tempering
We turn our attention to optimization. Critical scaling dynamics are intriguing from a physics perspective, but it is not essential that c-bits obey the same physics as p-bits in an optimization context. For optimization, we use an adaptive version of the parallel tempering algorithm, abbreviated as APT [55, 56, 57]. A given p-bit network is duplicated into $N$ system replicas, with each replica at a different inverse temperature $\beta$ (FIG. 1f). In fixed time intervals (time steps per swap attempt), a probabilistic swap attempt is conducted between adjacent replica pairs according to the Metropolis criterion:
@@
The APT algorithm can be applied using either p-bits or c-bits. While the neuron update rule is stochastic for p-bits (Eq. 1), it is deterministic for c-bits (Eq. 5). Replica swaps via the Metropolis criterion are always probabilistic (Eq. 15). Therefore, while p-bit APT is a purely probabilistic algorithm, c-bit APT is a hybrid chaotic-probabilistic scheme, where neuron updates are deterministic and replica swap attempts are probabilistic. In typical p-bit APT, the great majority of the random numbers are used for neuron updates rather than replica swap attempts. As such, if c-bit APT proves to be viable, then it would require orders of magnitude fewer random numbers. However, this potential benefit critically depends on the hardware cost of implementing c-bits, which may be greater than that of p-bits. Given the relative ease of creating pseudo random number generators through compact linear-feedback-shift-register circuits [58], we suspect that the c-bits may surreptitiously contain PRNG-like circuits in their implementation. Our view is that the p-bits and c-bits are very similar in their computational power, but p-bits enjoy a mathematical certainty that they sample from the Boltzmann distribution at steady state. We further discuss these ideas in Section VII.
The algorithm is referred to as ‘adaptive’ because we employ an instance-specific pre-processing method that finds a suitable $\beta$ schedule for a given problem. As shown in FIG. 5d, this process ensures that replica-swap probabilities are nearly uniform, avoiding bottlenecks during the exchange process (see Appendix .2). The complete APT algorithm is summarized in pseudo code (Algorithm 1).
@@ V.   Optimization of the 3D Spin Glass Problem
We believe that the hybrid chaotic-probabilistic APT algorithm we propose here is widely applicable to a whole host of other phase-based oscillator Ising machines. However, the simplicity of the c-bit and its state-based (rather than phase) representation through explicit latches may still be more appealing.
@@ VI.   Adaptive Parallel Tempering for Semiprime Factorization
Fig. 5: (a) p-bits are utilized to construct AND gates and full adders, which are the fundamental components for a multiplier. The product bits are clamped to a semiprime number. The system’s ground state finds the factors via invertible boolean logic [2]. (b) Energy profile of the 5 coldest replicas during hybrid chaotic-probabilistic APT. $E-E_{0}$ represents distance from ground state energy. (c) Time-to-solution vs. product size using a 10-bit $\times$ 10-bit factorizer for semiprime numbers. p-bit SA yields a slope of 0.441, while p-bit APT yields a slope of 0.438. c-bit SA exhibits a slope of 0.658, and hybrid APT has a slope of 0.420. (d) Average replica swap probabilities from factoring a specific semiprime number over 100 trials.
@@
where $Nt_{a}$ is the total computational time summed across all replicas. The time-to-solution represents the amount of computation necessary to factor semiprimes with a 99% confidence level (see Appendix .3). For a given product size (number of bits), we consider 180 unique semiprime numbers. We attempt to factor each semiprime number 80 times. $p(Nt_{a})$ is measured as the number of successful solves divided by the total number of attempted factorizations.
We devise a similar scheme for simulated annealing for the sake of comparison. We anneal 17 system replicas in an embarrassingly parallel manner. Each replica runs for $t_{a}=10^{5}$ time steps, and $\beta$ is linearly increased from 0 to 10. In order to avoid random-find effects, we only check for the correct factors at the last 1000 time steps. If any of the 17 replicas finds the correct factors, the trial is considered a successful solve. For a given problem size (# bits), we again consider 180 unique semiprimes. We attempt to factor each semiprime 120 times. $p(Nt_{a})$ and time-to-solution are calculated in a fashion analogous to APT.
@@
Using this method of comparison, FIG. 5c shows that purely probabilistic APT exhibits remarkably similar algorithmic scaling to the newly introduced chaotic-probabilistic hybrid scheme. Moreover, they both exhibit better scaling than SA-based approaches. Pure chaotic SA, which has been commonly implemented [38], shows the worst scaling. Furthermore, both p-bits and c-bits observe a boundary effect as the problem size approaches 20 bits, where the time-to-solution actually decreases. This is a feature particular to the fixed 20-bit factorizer circuit: since numbers larger than 20 bits cannot be represented by the circuit, the solution space is reduced as we approach 20 bits. We exclude this data because it is uninformative for algorithmic scaling.
The two computing schemes not only share similar algorithmic scaling but also exhibit similarities in their underlying mechanisms. We consider a specific 11-bit semiprime number and conduct APT for $t_{a}=10^{5}$ time steps. We average over 100 random seeds. FIG. 5d shows the probability of the Metropolis criterion dictating a swap between a given pair of replicas. Remarkably, the hybrid APT scheme follows a similar qualitative distribution as the probabilistic APT scheme.
@@ VII.   Possible Hardware Realizations
Before we conclude, we make a few remarks regarding hardware implementations of chaotic networks. Chaotic bits have typically been implemented in digital and analog CMOS [45, 44]. Similar to CMOS implementations of p-bits, the main appeal of c-bit networks is in their asynchronous and massively parallel operation, provided that the synaptic input (Eq. 2) and the latching mechanism can be computed quickly to avoid detrimental parallel updates. We suspect that the initial randomization of oscillator phases and round-off errors in floating point software simulation are largely responsible for adding a layer of stochasticity to these networks. In hardware implementation, the needed noise can come from thermal noise in the environment. Beyond digital implementation, c-bits can make use of “natural” latches. For example, the physics of spin-torque switching of a non-volatile nanomagnet [62] can function as the latch: the magnet will switch only if a current threshold is reached, and the magnet will exhibit hysteresis until the current threshold changes sign, much like the billiard-ball metaphor used in c-bits.
A purported benefit of c-bits is the absence of explicit random numbers in their update rules. However, this claim requires qualifications. First, low-cost pseudo random number generators, such as LFSRs, do not significantly increase the complexity of implementing a p-bit. For example, an $n$-bit LFSR requires about $32\times n$ transistors to implement [58]. In contrast, constructing a c-bit digitally may be significantly more complex than this cost due to the necessity of an oscillator with a tunable duty cycle. Additionally, there may be “hidden” sources of randomness, such as floating point round-off errors in simulation or thermal noise in physical implementation.
Augmenting c-bits with explicit randomness in powerful tempering algorithms such as the one considered in this paper could offer further benefits. We believe that while the lack of random numbers may be an important c-bit advantage, comparisons must be made in concrete implementations in order to be meaningful (for example, against the most energy-efficient and compact hardware solutions of p-bits, using magnetic nanodevices [3]). Nevertheless, we believe that the similar performance of c-bits and p-bits across the wide variety of problems considered in this paper is very promising, providing valuable insights into the level of stochasticity required for effective probabilistic computing.
@@ Conclusion
In this work, we evaluated and improved deterministic chaotic bit networks that combine billiard-like oscillators with latches. We demonstrated that augmenting these purely deterministic devices with stochasticity exhibits algorithmic scaling comparable to fully stochastic networks of p-bits. Additionally, we achieved better algorithmic scaling than common simulated annealing-based approaches for c-bits. Notably, the proposed algorithm is applicable to any other oscillator Ising machine.
Moreover, we have shown that c-bits approximately sample from the quantum Boltzmann law in a 1D TFIM model. Although c-bits exhibited qualitatively different critical scaling in the 2D Ising model and the 3D spin glass problem compared to the exact probabilistic bit formulations, their use in practical optimization problems remains promising. Combining stochastic p-bits and deterministic c-bits in asynchronous and massively parallel hardware implementations could create powerful domain-specific computers for combinatorial optimization and probabilistic sampling.
Acknowledgements
We gratefully acknowledge Sanaaya Lakdawala and Navid Anjum Aadit for useful initial discussions. We thank Nikhil Shukla for insights regarding Oscillator Ising Machines and Masoud Mohseni for useful discussions on the Kibble-Zurek mechanism. This work is partially supported by an Office of Naval Research Young Investigator Program grant, and a National Science Foundation CCF 2106260 grant. Use was made of computational facilities purchased with funds from the National Science Foundation (CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (MRSEC; NSF DMR 2308708) at UC Santa Barbara.
@@ Methods
As the system is annealed towards the critical point, the equilibration time increases. If the initial $\lambda$ is far away from the critical point, the system quickly relaxes to its equilibrium and can follow the annealing schedule in quasi-equilibrium or adiabatically. There comes a point, $\hat{t}$, when the remaining annealing time, $t_{a}-\hat{t}$ is precisely equal to the relaxation time of the system, $\tau[\lambda(\hat{t})]$. At this point, any further change in $\lambda$ that increases the relaxation time results in the system not being able to equilibrate by the end of the annealing. The idea of KZM, then, is to find this “freeze-out” boundary $\hat{t}$ and to relate non-equilibrium correlation lenghts to equilibrium correlations lengths via $\xi[\lambda(\hat{t})]$, which is then related to a density of topological defects via dimensional analysis. Based on this sketch, we proceed to find $\hat{t}$ by equating $|t_{a}-\hat{t}|$ to $\tau(\hat{\lambda})$ using Eq. 19 and Eq. 20. We obtain:

### v2

@@ IV.   Adaptive Parallel Tempering
Critical scaling dynamics are intriguing from a physics perspective, but it is not essential that c-bits obey the same physics as p-bits in an optimization context. For optimization, we use an adaptive version of the parallel tempering algorithm, abbreviated as APT [60, 61, 62]. A given p-bit or c-bit network is duplicated into $N$ system replicas, with each replica at a different inverse temperature $\beta$ (FIG. 1f). In fixed time intervals (time steps per swap attempt), a probabilistic swap attempt is conducted between adjacent replica pairs according to the Metropolis criterion:
@@
The APT algorithm can be applied using either p-bits or c-bits. While the neuron update rule is stochastic for p-bits (Eq. 1), it is deterministic for c-bits (Eq. 5). Replica swaps via the Metropolis criterion are always probabilistic (Eq. 21). Therefore, while p-bit APT is a purely probabilistic algorithm, c-bit APT is a hybrid chaotic-probabilistic scheme, where neuron updates are deterministic and replica swap attempts are probabilistic. In typical p-bit APT, the great majority of the random numbers are used for neuron updates rather than replica swap attempts. In contrast, c-bit APT requires orders of magnitude fewer random numbers. However, this potential benefit critically depends on the hardware cost of implementing c-bits, which may be greater than that of p-bits. Given the relative ease of creating pseudo random number generators through compact linear-feedback-shift-register circuits [63], we suspect that the c-bits may surreptitiously contain PRNG-like circuits in their implementation. While p-bits and c-bits appear to be similar in their computational power, it is important to note that only p-bits enjoy a mathematical certainty that they sample from the Boltzmann distribution at steady state. We further discuss these ideas in Section VII.
The algorithm is referred to as ‘adaptive’ because we employ an instance-specific pre-processing method that finds a suitable $\beta$ schedule for a given problem. As shown in FIG. 6e, this process ensures that replica-swap probabilities are nearly uniform, avoiding bottlenecks during the exchange process (see Appendix .2). The complete APT algorithm is summarized in pseudo code (Algorithm 1).
@@ V.   Optimization of the 3D Spin Glass Problem
The hybrid chaotic-probabilistic APT algorithm we propose here may be applicable to other approaches, such as the noise-injected oscillator-based Ising machines proposed in Ref. [33], that are designed to enable sampling from the Boltzmann distribution. However, the simplicity of the c-bit and its state-based (rather than phase) representation through explicit latches may still be more appealing.
@@ VI.   Adaptive Parallel Tempering for Semiprime Factorization
Fig. 6: (a) p-bits are utilized to construct AND gates and full adders, that are the fundamental gates for a multiplier. The product bits are clamped to a semiprime number as in invertible boolean logic [2]. (b) Energy profile of the 5 coldest replicas during hybrid chaotic-probabilistic APT. $E-E_{0}$ represents distance from ground state energy. (c) $\beta$ profile obtained from APT preprocessing (see Appendix .2). (d) Time-to-solution vs. product size using a 10-bit $\times$ 10-bit factorizer for semiprime numbers. p-bit SA yields a slope of 0.441, while p-bit APT yields a slope of 0.438. c-bit SA exhibits a slope of 0.658, and hybrid APT has a slope of 0.420. (e) Average replica swap probabilities from factoring a specific semiprime number over 100 trials.
@@
where $Nt_{a}$ is the total computational time summed across all replicas. The time-to-solution represents the amount of computation necessary to factor semiprimes with a 99% success probability (see Appendix .3). For a given product size (number of bits), we consider 180 unique semiprime numbers. We attempt to factor each semiprime number 80 times. $p(Nt_{a})$ is measured as the number of successful solves divided by the total number of attempted factorizations.
We devise a similar scheme for simulated annealing for the sake of comparison. We anneal 17 system replicas in parallel. Each replica runs for $t_{a}=10^{5}$ time steps, and $\beta$ is linearly increased from 0 to 10. In order to avoid random-find effects due to small problem sizes, we only check for the correct factors at the last 1000 time steps. If any of the 17 replicas finds the correct factors, the trial is considered a successful solve. For a given problem size (# bits), we again consider 180 unique semiprimes. We attempt to factor each semiprime 120 times. $p(Nt_{a})$ and time-to-solution are calculated in a fashion analogous to APT.
@@
Using this method of comparison, FIG. 6d shows that purely probabilistic APT exhibits remarkably similar algorithmic scaling to the newly introduced chaotic-probabilistic hybrid scheme. Moreover, they both exhibit better scaling than SA-based approaches. Pure chaotic SA, which has been commonly implemented [40], shows the worst scaling. Furthermore, both p-bits and c-bits observe a boundary effect as the problem size approaches 20 bits, where the time-to-solution actually decreases (for all the algorithms). This is a feature particular to the fixed 20-bit factorizer circuit: since numbers larger than 20 bits cannot be represented by the circuit, the solution space is reduced as we approach 20 bits. We exclude this data because it is uninformative for algorithmic scaling.
The two computing schemes not only share similar algorithmic scaling but also exhibit similarities in their underlying mechanisms. We consider a specific 11-bit semiprime number and conduct APT for $t_{a}=10^{5}$ time steps. We average over 100 random seeds. FIG. 6e shows the probability of the Metropolis criterion dictating a swap between a given pair of replicas. Remarkably, the hybrid APT scheme follows a similar qualitative distribution as the probabilistic APT scheme.
@@ VII.   Possible Hardware Realizations
Based on the original c-bit definition that is different the one we used in this paper, chaotic bit implementations have been made in digital and analog CMOS [51, 50, 67]. We now make some remarks regarding hardware implementations of chaotic networks, however, a concrete discussion of hardware implementations is beyond our scope.
As a first evaluation of the algorithm we proposed, the results we presented are hardware-agnostic based on scaling laws and algorithmic exponents. However, there are restrictions that would be encountered in any hardware implementation.
The first point is related to the synapse calculation time. In our software implementation, the synapse (Eq. 2) is updated every time step, so c-bits states are always updated using up-to-date information from their neighbors. Just as in asynchronous networks of p-bits [45], the asynchronous operation of c-bits require a fast local field calculation. This places stringent requirements on the network topology since dense networks will require more time to update the synapse.
The second point is related to the latch mechanism: as we showed in FIG. 2, if the phases of two different billiards are too close to each in other in time, the latches might erroneously update at the same time depending on the way the latch is implemented, especially in dense network topologies.
Beyond digital implementations, c-bits can make use of “natural” latches. For example, the physics of spin-torque switching of a non-volatile nanomagnet [68] can function as the latch: the magnet will switch only if a current threshold is reached, and the magnet will exhibit hysteresis until the current threshold changes sign, much like the billiard-ball metaphor used in c-bits.
A purported benefit of c-bits is the absence of explicit random numbers in their update rules. However, it is not clear whether the cost of c-bit hardware implementation is significantly cheaper than that of p-bits. First, low-cost pseudo random number generators, such as LFSRs, do not significantly increase the complexity of implementing a p-bit. For example, an $n$-bit LFSR requires about $32\times n$ transistors to implement [63]. In contrast, constructing a c-bit digitally may be significantly more complex than this cost due to the necessity of an oscillator with a tunable duty cycle. Additionally, there may be “hidden” sources of randomness, such as floating point round-off errors in simulation or thermal noise in physical implementation.
Augmenting c-bits with explicit randomness in powerful tempering algorithms such as the one considered in this paper could offer further benefits. We believe that while the lack of random numbers may be an important c-bit advantage, comparisons must be made in concrete implementations in order to be meaningful (for example, against the most energy-efficient and compact hardware solutions of p-bits, using magnetic nanodevices [3]). We refrain from making preemptive assumptions about the scalability and efficiency of c-bit hardware, as specific implementation details are unknown. p-bit Monte Carlo sweeps are not 1-to-1 with c-bit’s time variable $t$, so speculations are not particularly informative until implementation. Nevertheless, we believe that the similar algorithmic scaling performance of c-bits and p-bits across the wide variety of problems considered in this paper is very promising, providing valuable insights into the level of stochasticity required for effective probabilistic computing.
@@ Conclusion
In this work, we evaluated and improved deterministic chaotic bit networks that combine billiard dynamics with latches. We note that while algorithm parameters were not excessively tuned, using similar parameters for p-bit and c-bit versions of SA and APT enabled fair comparison of algorithmic scaling performance. We demonstrated that augmenting deterministic c-bits with stochasticity via probabilistic APT replica swaps exhibits algorithmic scaling comparable to APT using fully stochastic networks of p-bits. Additionally, noise-injected c-bit APT achieved better algorithmic scaling than common simulated annealing-based algorithms. Notably, the proposed adaptive parallel tempering algorithm is applicable to certain classes of oscillator Ising machines and similar nonlinear devices.
Moreover, we have shown that c-bits approximately sample from the quantum Boltzmann law in a 1D TFIM model. On the 2D Ising and 3D spin glass problems, c-bits qualitatively exhibit similar critical scaling dynamics as p-bits, though a precise correspondence between their measured scaling exponents $\kappa_{c}$ is unclear. While intriguing from a physics perspective, these critical dynamics are not essential to optimization, for which c-bits show strong promise. Combining stochastic p-bits and deterministic c-bits in asynchronous and massively parallel hardware implementations could create powerful domain-specific computers for combinatorial optimization and probabilistic sampling.
@@ Methods
As the system is annealed towards the critical point, the equilibration time increases. If the initial $\lambda$ is far away from the critical point, the system quickly relaxes to its equilibrium and can follow the annealing schedule in quasi-equilibrium or adiabatically. There comes a point, $\hat{t}$, when the remaining annealing time, $t_{a}-\hat{t}$ is precisely equal to the relaxation time of the system, $\tau[\lambda(\hat{t})]$. At this point, any further change in $\lambda$ that increases the relaxation time results in the system not being able to equilibrate by the end of the annealing. The idea of KZM, then, is to find this “freeze-out” boundary $\hat{t}$ and to relate non-equilibrium correlation lengths to equilibrium correlations lengths via $\xi[\lambda(\hat{t})]$, which is then related to a density of topological defects via dimensional analysis. Based on this sketch, we proceed to find $\hat{t}$ by equating $|t_{a}-\hat{t}|$ to $\tau(\hat{\lambda})$ using Eq. 25 and Eq. 26. We obtain:
@@
Acknowledgements
We gratefully acknowledge discussions with Sanaaya Lakdawala and Navid Anjum Aadit, Nikhil Shukla and Corentin Delacour for insights regarding Oscillator Ising Machines and Masoud Mohseni for insights on KZM. This work is partially supported by an Office of Naval Research Young Investigator Program grant, and a National Science Foundation CCF 2106260 grant. Use was made of computational facilities purchased with funds from the National Science Foundation (CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (MRSEC; NSF DMR 2308708) at UC Santa Barbara.